# Booster: Tackling Harmful Fine-tuning for Large Language Models via Attenuating Harmful Perturbation

### v3

@@ Abstract
Harmful fine-tuning issue (Qi et al., 2023) poses serious safety concerns for Large language models’ fine-tuning-as-a-service. While existing defenses have been proposed to mitigate the issue, their performances are still far away from satisfactory, and the root cause of the problem has not been fully recovered. To this end, we in this paper show that harmful perturbation over the model weights could be a cause of alignment-broken of harmful fine-tuning. In order to attenuate the negative impact of harmful perturbation, we propose an alignment-stage solution, dubbed Booster. Technically, along with the original alignment loss, we append a loss regularizer in the alignment stage’s optimization. The regularizer ensures that the model’s harmful loss reduction before/after simulated harmful perturbation is attenuated, thereby mitigating the subsequent fine-tuning risk. Empirical results show that Booster can effectively reduce the harmful score of the fine-tuned models while maintaining the performance of downstream tasks. Our code is available at https://github.com/git-disl/Booster.
@@ 1 Introduction
Fine-tuning-as-a-service has been a successful business service model adopted by main-stream LLM service providers  [^1] Fine-tuning API by OpenAI: https://platform.openai.com/docs/guides/fine-tuning.. It aims to deliver customized LLM service to users by asking them to upload demonstration data using the desired pattern. Then the service provider will fine-tune the foundation models on users’ behalf. However, recent studies (Qi et al., 2023; Yang et al., 2023; Zhan et al., 2023) uncovered a harmful fine-tuning issue, which shows with red-teaming that a few harmful data contained in the user fine-tuning dataset can trigger the fine-tuned models to forget the safety alignment enforced before. This vulnerability renders a large attack surface, downgrading the service quality and sustainability, and there is still no solution with great scalability to tackle the problem.
Existing solutions against harmful fine-tuning issues can be classified into three main categories according to which stage the mitigation is introduced, i.e., i) alignment-stage solution (Huang et al., 2024c; Rosati et al., 2024a; Tamirisa et al., 2024), ii) fine-tuning-stage solution(Mukhoti et al., 2023; Huang et al., 2024b), and iii) post-fine-tuning stage solution (Yi et al., 2024; Hsu et al., 2024; Huang et al., 2024a). Among the three categories, alignment-stage solutions draw the broadest interest thanks to their computation efficiency. Specifically, mitigation solutions in most cases come with extra computation, which means for fine-tuning-stage solutions and post-fine-tuning stage solutions, extra computation comes with every incoming fine-tuning request. On the contrary, only one-time overhead is required for an alignment-stage solution. Due to its superiority, we in this paper aim to promote the alignment-stage solution.
Vaccine (Huang et al., 2024c) and RepNoise (Rosati et al., 2024a) are two representative alignment-stage solutions. (Huang et al., 2024c) propose the concept of harmful embedding drift, which measures the drift of embedding over the alignment data before and after fine-tuning. The proposed method named Vaccine solves a mini-max solution to mitigate the impact of the embedding drift. However, Vaccine solely uses the alignment dataset (i.e., harmful prompt-safe answer pair), which alone may not be sufficient enough to counter the harmful attack. (Rosati et al., 2024a) further introduce RepNoise that utilizes the harmful dataset (harmful prompt-harmful answer pair). The idea is to introduce an MMD regularizer to perturb the embedding of the harmful data such that its distribution reduces to a normalized Gaussian noise. However, RepNoise may fall short given that harmful fine-tuning is still able to reshape the harmful embedding distribution under some conditions.
@@
Driven by this research question, we in this paper first study how the harmful data is going to reshape the aligned model. We discover that harmful perturbation, i.e., taking optimization direction over the harmful data to the model, contributes to the reduction of the harmful loss and therefore should be the culprit of the alignment-broken effect. To attenuate its negative impact, we propose a regularizer to utilize the simulated harmful dataset to reduce its harmful loss reduction rate. Combining the regularizer with the original alignment loss over the alignment data, we form the optimization problem, which is then efficiently solved by an iterative gradient method named Booster. Empirical results show that Booster outperforms existing solutions by respectively reducing up-to 17.26% and 20.08% average harmful score compared with Vaccine and RepNoise, while maintaining the same level of fine-tune accuracy.
@@ 2 Related Work
Safety Alignment. Large language model’s safety alignment concerns on how to regularize the model’s output such that the model is able to output a refusal answer whenever a harmful prompt is given. The mainstream techniques to achieve alignment includes supervised fine-tuning (SFT) and RLHF (Ouyang et al., 2022; Dai et al., 2023; Bai et al., 2022; Wu et al., 2023; Dong et al., 2023; Rafailov et al., 2023; Yuan et al., 2023). Both techniques exploit a safety alignment dataset, which consists of demonstration data showing how to give refusal answers to harmful prompts. The pre-trained models are then trained on these data with SFT or RLHF to achieve safety alignment.
@@
Harmful Fine-tuning. However, recent studies (Qi et al., 2023; Yang et al., 2023; Zhan et al., 2023) show that models aligned by SFT or RLHF can be jail-broken after fine-tuning on a partially harmful dataset, i.e., the model forgets to give refusal answer towards harmful prompts after fine-tuning on a few harmful samples. Efforts have been made to analyze the mechanism of the attack. Huang et al. (2024c) accounts for the reason of forgetting as harmful embedding drift. Leong et al. (2024) discuss the attack mechanisms over two different attack settings, and Peng et al. (2024) propose a safety metric to measure the attack impact over different models. Existing mitigation solutions to the issue can be classified into three categories according to which stage the mitgation is launched, i.e., alignment-stage solution (Huang et al., 2024c; Rosati et al., 2024a; b), fine-tuning-stage solution (Mukhoti et al., 2023; Huang et al., 2024b; Lyu et al., 2024; Wang et al., 2024; Qi et al., 2024), and post-fine-tuning stage solution (Hsu et al., 2024; Yi et al., 2024; Huang et al., 2024a). The method proposed in this paper should be classified into an alignment-stage solution, whose objective is to vaccinate the LLM such that it can be more robust to the attack launched after the alignment.
@@
Meta Learning. Meta learning (Finn et al., 2017; Rajeswaran et al., 2019) is a technique aiming to produce a meta-model that can be generalized to different downstream tasks before fine-tuning occurs. In order to achieve this goal, the optimization objective is over the model after a gradient step, i.e., $\sum_{i\in\mathcal{T}}f_{i}(w-\nabla f(w))$ where $\mathcal{T}$ is a set of downstream tasks. In this paper, the proposed booster method also involves one-step gradient information into the optimization objective, aiming to attenuate the impact of harmful perturbation before fine-tuning occurs.
@@
To the best of our knowledge, we are the first to identify harmful perturbation as the cause of alignment broken, and this is our first solution that utilizes this concept to design a defense to strengthen the model’s robustness. The defense can be combined with other existing defense solutions, e.g., Vaccine (Huang et al., 2024c) and RepNoise (Rosati et al., 2024a) to further improve defense performance. A concurrent research TAR (Tamirisa et al., 2024) utilizes a similar meta-learning technique to simulate the harmful perturbation in the alignment stage. However, the insight as well as the design of our method is different from theirs. See Appendix C for a detailed discussion.
@@ 3.1 Harmful fine-tuning
Considered Scenario. Harmful fine-tuning is a security issue faced by the LLM service provider (e.g., OpenAI). The considered scenario is that users upload a few pieces of data (a subset of them are harmful) to the service provider, and the service provider fine-tune their safety-aligned foundation model with those provided datasets. The fine-tuned data are deployed in the service provider’s server and are used to serve personalized output to the users.
Threat models . We assume that the users upload a user fine-tuning dataset with $p$ (percentage) of harmful data, and other $1-p$ (percentage) of data are benign fine-tuning data. The harmful data and the benign data are considered inseparable (Qi et al., 2023; Huang et al., 2024c; Rosati et al., 2024a).
Assumptions. We assume the service provider maintains an alignment dataset (harmful prompt-safe answer pairs) and a harmful dataset (harmful prompt-harmful answer pairs) for alignment. The availability of the two pairs of data is also made in (Rosati et al., 2024a; Tamirisa et al., 2024) and both the two pairs of data are available in existing open datasets (e.g., BeaverTail (Ji et al., 2023)).
@@ 3.2 Harmful Perturbation
Next, we try to explore the concept of harmful perturbation to show the reason for alignment failure and motivate the design of our proposed defense.
@@
Definition of harmful perturbation. We refer to harmful perturbation as taking one step towards the gradient direction over the harmful data. In the fine-tuning stage, the model update rule will be $\bm{w}_{t+1}=\bm{w}_{t}-\eta g(\bm{w}_{t})$ where $g(\bm{w}_{t})$ is a stochastic gradient over a random sample of user data. If $g(\bm{w}_{t})$ is a stochastic gradient from a harmful data, and we take this stochastic gradient over $\bm{w}_{t}$, this results in the so-called harmful perturbation, which should be the root reason leading to a successful harmful fine-tuning attack. We next utilize experimental results to demonstrate the impact of harmful perturbation, and how it leads to a successful attack.
@@
Impact of harmful perturbation. The left of Figure 2 shows that the model’s harmful score will substantially increase along with optimization steps invested in fine-tuning on a pure harmful dataset. On the contrary, the harmful score will not be affected much via fine-tuning on a pure SST2 dataset. This indicates that taking a step with the gradient of the harmful data (i.e., harmful perturbation) is indeed the reason for the alignment broken. We next show in the middle of Figure 2 how taking harmful perturbation and SST2 perturbation in each epoch will affect the harmful training loss. As shown, harmful perturbation significantly reduces the harmful training loss, which means the model starts to fit the harmful data. On the contrary, training on SST2 would only slightly increase the harmful training loss. The right of Figure 2 shows a similar trend, which indicates that the fitting to the training harmful data can generalize to other unseen harmful data. In summary, our finding concludes that harmful perturbation (taking a step over the gradient of the harmful data) contributes to the reduction of harmful training/testing loss, and eventually makes the model respond in a harmful way.
@@
Figure 2: Model statistics (Left: harmful score, Middle: harmful training loss, Right: harmful testing loss) after fine-tuning on pure SST2/harmful data for different steps. Specially, harmful training loss refers to the loss over the harmful data used in training, while harmful testing loss refers to that over the testing harmful data which the model never see in fine-tuning phase.
Derived Insight. Because harmful fine-tuning data is considered to be inseparable from the benign fine-tuning data (e.g., SST2 ) (Huang et al., 2024c; Rosati et al., 2024a), harmful perturbation is indeed inevitable in the user fine-tuning stage. Given this, it is essential to increase the model’s resilience towards the harmful perturbation in the earlier stage, i.e., alignment stage, which however, is not sufficiently explored by the existing literature.
@@ 4 Methodology
To mitigate the harmful perturbation issue, we propose an alignment stage solution to attenuate the impact of the harmful perturbation that will inevitably be incurred in the fine-tuning stage.
Explicitly, we want to solve the following optimization problem in the alignment stage:
@@
where $f(\bm{w})$ is the empirical loss  [^2] The empirical loss here is cross-entropy loss if we use SFT for alignment/fine-tuning. The loss can be more complicated if RLHF is adopted. over the alignment dataset and $h(\bm{w})$ is the empirical loss over the harmful dataset, $\lambda$ is the regularizer’s intensity, and $\alpha$ is the step size. Our contribution lies in the second term, which measures the gap between the original harmful loss and the harmful loss after taking a normalized step with the harmful gradient. The idea is to minimize the impact of potential harmful perturbation towards the alignment model while simultaneously minimizing its alignment loss. Specifically, $h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})$ simulates the loss after one normalized step of fine-tuning on harmful samples, and $h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})$ simulates the decrease of harmful loss after one step of fine-tuning on harmful samples. By minimizing this gap, the decrease of harmful loss after taking optimization on the real harmful samples in the fine-tuning stage will be minimized (i.e., impact of harmful perturbation will be attenuated).
@@
where $\eta$ is the learning rate. Note that the term $\nabla(\bm{w}_{t}-\alpha\frac{\nabla h(\bm{w}_{t})}{\|\nabla h(\bm{w}_{t})\|})$ contains second order information (i.e., Hessian Matrix), which is very computation expensive to obtain. Inspired by (Finn et al., 2017; Rajeswaran et al., 2019), we rewrite the update rule by approximating this second-order gradient term to be constant, as follows:
@@
0:  Regularizer intensity, $\lambda$; Step size, $\alpha$; Learning rate, $\eta$;
@@
We present the detailed algorithm in Algorithm 1. The overall procedure requires three forward/backward passes for each optimization step. The first pass is to evaluate the gradient over a batch of alignment data $(\bm{x}_{t},\bm{y}_{t})$ and obtain ~$\tilde{\nabla}f(\bm{w}_{t})$. The second pass is to evaluate the harmful gradient over a batch of harmful data and obtain ~$\tilde{\nabla}h(\bm{w}_{t})$, and the third pass is to evaluate the harmful gradient again after taking a normalized harmful gradient step. After collecting all the three gradient components, we take the final gradient step by Eq. (3).
The proposed algorithm is named Booster, named after another alignment stage solution "Vaccine" (Huang et al., 2024c). Of note, the design idea of Booster are completely different from Vaccine. In a high level, Booster is concerned on how to utilize harmful dataset to simulate the harmful perturbation, while Vaccine is concerned on how to mitigate the hidden embedding drift over the alignment dataset. Factualy, the two techniques can be combined and we will discuss in Appendix D.

### v4

@@ Abstract
Harmful fine-tuning attack poses serious safety concerns for large language models’ fine-tuning-as-a-service. While existing defenses have been proposed to mitigate the issue, their performances are still far away from satisfactory, and the root cause of the problem has not been fully recovered. To this end, we in this paper show that harmful perturbation over the model weights could be a probable cause of alignment-broken. In order to attenuate the negative impact of harmful perturbation, we propose an alignment-stage solution, dubbed Booster. Technically, along with the original alignment loss, we append a loss regularizer in the alignment stage’s optimization. The regularizer ensures that the model’s harmful loss reduction after the simulated harmful perturbation is attenuated, thereby mitigating the subsequent fine-tuning risk. Empirical results show that Booster can effectively reduce the harmful score of the fine-tuned models while maintaining the performance of downstream tasks. Our code is available at https://github.com/git-disl/Booster.
Disclaimer: This document contains content that some may find disturbing or offensive, including content that is hateful or violent in nature.
@@ 1 Introduction
Fine-tuning-as-a-service has been a successful business service model adopted by main-stream Large Language model (LLM) service providers  [^1] Fine-tuning API by OpenAI: https://platform.openai.com/docs/guides/fine-tuning.. It aims to deliver customized LLM service to users by asking them to upload demonstration data using the desired pattern. Then the service provider will fine-tune the foundation models on users’ behalf. However, recent studies (Qi et al., 2023; Yang et al., 2023; Zhan et al., 2023; Lermen et al., 2023; Yi et al., 2024a) uncovered a harmful fine-tuning issue, which shows with red-teaming that a few harmful data contained in the user fine-tuning dataset can trigger the fine-tuned models to forget the safety alignment enforced before. This vulnerability renders a large attack surface, downgrading the service quality and sustainability, and there is still no solution with great scalability to tackle the problem.
Existing solutions against harmful fine-tuning issues can be classified into three main categories according to which stage the mitigation is introduced, i.e., i) alignment-stage solution (Huang et al., 2024e; Rosati et al., 2024b; Tamirisa et al., 2024), ii) fine-tuning-stage solution(Mukhoti et al., 2023; Huang et al., 2024d), and iii) post-fine-tuning stage solution (Yi et al., 2024c; Hsu et al., 2024; Huang et al., 2024a). Among the three categories, alignment-stage solutions draw the broadest interest thanks to their computation efficiency. Specifically, mitigation solutions in most cases come with extra computation, which means for fine-tuning-stage solutions and post-fine-tuning stage solutions, extra computation comes with every incoming fine-tuning request. On the contrary, only one-time overhead is required for an alignment-stage solution. Due to its superiority, we in this paper aim to promote the alignment-stage solution.
Vaccine (Huang et al., 2024e) and RepNoise (Rosati et al., 2024b) are two representative alignment-stage solutions. (Huang et al., 2024e) propose the concept of harmful embedding drift, which measures the drift of embedding over the alignment data before and after fine-tuning. The proposed method named Vaccine solves a mini-max solution to mitigate the impact of the embedding drift. However, Vaccine solely uses the alignment dataset (i.e., harmful prompt-safe answer pair), which alone may not be sufficient enough to counter the harmful attack. (Rosati et al., 2024b) further introduces RepNoise that utilizes the harmful dataset (harmful prompt-harmful answer pair). The idea is to introduce an MMD regularizer to perturb the embedding of the harmful data such that its distribution reduces to a normalized Gaussian noise. However, RepNoise may fall short given that harmful fine-tuning is still able to reshape the harmful embedding distribution under some conditions.
@@
Driven by this research question, we in this paper first study how the harmful data is going to reshape the aligned model. We discover that harmful perturbation, i.e., taking optimization direction over the harmful data to the model, contributes to the reduction of the harmful loss and therefore should be the culprit of the alignment-broken effect. To attenuate its negative impact, we propose a regularizer to utilize the simulated harmful dataset to reduce the harmful loss reduction rate after harmful perturbation. Combining the regularizer with the original alignment loss over the alignment data, we form the optimization problem, which is then efficiently solved by an iterative gradient method named Booster. Empirical results show that Booster outperforms existing solutions by respectively reducing up-to 17.26% and 20.08% average harmful scores compared with Vaccine and RepNoise, while maintaining the same level of fine-tune accuracy.
@@ 2 Related Work
Safety Alignment. Large language model’s safety alignment concerns how to regularize the model’s output such that the model is able to output a refusal answer whenever a harmful prompt is given. The mainstream techniques to achieve alignment includes supervised fine-tuning (SFT), RLHF (Ouyang et al., 2022; Dai et al., 2023; Bai et al., 2022; Wu et al., 2023; Dong et al., 2023; Rafailov et al., 2023; Yuan et al., 2023), and a few other research, e.g., Stable Alignment (Liu et al., 2023), Selfee (Ye et al., 2023), Circuit Breakers (Zou et al., 2024), 3HOptimization Yang et al. (2025) and $H^{3}$usion (Tekin et al., 2024). Both SFT and RLHF exploit a safety alignment dataset, which consists of demonstration data showing how to give refusal answers to harmful prompts. The pre-trained models are then trained on these data with SFT or RLHF to achieve safety alignment.
@@
Harmful Fine-tuning Attack. However, recent studies on harmful fine-tuning attack (Qi et al., 2023; Yang et al., 2023; Zhan et al., 2023; Lermen et al., 2023; Yi et al., 2024a) show that models aligned by SFT or RLHF can be jail-broken after fine-tuning on a partially harmful dataset, i.e., the model forgets to give refusal answer towards harmful prompts after fine-tuning on a few harmful samples. More advanced attacks have been proposed in (He et al., 2024; Halawi et al., 2024; Poppi et al., 2024; Hawkins et al., 2024; Xu et al., 2025; Huang et al., 2025). Efforts have been made to analyze the mechanism of the attack. (Huang et al., 2024e) accounts for the reason of forgetting as harmful embedding drift. (Leong et al., 2024) discuss the attack mechanisms over two different attack settings, and (Peng et al., 2024) propose a safety metric to measure the attack impact over different models. (Hsiung et al., 2025) gives some empirical insights of alignment data selection, and (Guo et al., 2024) analyze the attack mechanism for vision LLM. Existing mitigation solutions to the issue can be classified into three categories according to which stage the mitigation is launched: i) alignment-stage solution, including Vaccine (Huang et al., 2024e), RepNoise(Rosati et al., 2024b), CTRL (Liu et al., 2024c), TAR (Tamirisa et al., 2024), SN-Tune (Zhao et al., 2025), T-Vaccine (Liu et al., 2024a) and KT-IPA (Cheng et al., 2025), ii) fine-tuning-stage solution, including LDIFS (Mukhoti et al., 2023), Freeze (Wei et al., 2024), Constrain-SFT (Qi et al., 2024a), Paraphrase (Eiras et al., 2024), ML-LR (Du et al., 2024), Freeze+ (Li & Kim, 2025), SaLoRA (Li et al., 2025), SafeInstr (Bianchi et al., 2023), VLGuard (Zong et al., 2024), Lisa (Huang et al., 2024d), BEA(Wang et al., 2024a), PTST (Lyu et al., 2024), Seal (Shen et al., 2024), SAFT (Choi et al., 2024), SPPFT(Li et al., 2024a), CMRM(Liu et al., 2024b). iii) post-fine-tuning stage solution, including LAT (Casper et al., 2024), SOMF (Yi et al., 2024c), Safe LoRA (Hsu et al., 2024), Antidote (Huang et al., 2024a), SafetyLock (Zhu et al., 2024), IRR (Wu et al., 2024), NLSR (Yi et al., 2024b), LoRA fusion (Gudipudi et al., 2024), Panacea (Wang et al., 2025) and BEAT(Yi et al., 2025). Several study investigate harmful fine-tuning attack in other scenarios, e.g., federated learning(Ye et al., 2024; Li et al., 2024b) and diffusion models(Pan et al., ). For a more comprehensive literature review, we refer to surveys (Huang et al., 2024c; Reuel et al., 2024; Sicari et al., 2024; Barez et al., 2025; Verma et al., 2024; Cheng et al., 2024; Cui et al., 2024) for more discussion. The method proposed in this paper should be classified into an alignment-stage solution, whose objective is to vaccinate the LLM such that it can be more robust to the attack launched after the alignment. A recent study (Qi et al., 2024b; Rosati et al., 2024a) suggests that existing alignment stage defenses, i.e., TAR (Tamirisa et al., 2024) and RepNoise (Rosati et al., 2024b) fail in some corner attack cases, potentially exposing vulnerability of this category of defenses. With that said, we advocate that continuous research on alignment stage solution is still necessary, even though they might face a more unpredictable and sophisticated corner attack case under the minimal assumption made on the defender’s side.
@@
Meta Learning. Meta learning (Finn et al., 2017; Rajeswaran et al., 2019) is a technique aiming to produce a meta-model that can be generalized to different downstream tasks before fine-tuning occurs. In order to achieve this goal, the optimization objective is over the model after a gradient step, i.e., $\sum_{i\in\mathcal{T}}f_{i}(\bm{w}-\nabla f(\bm{w}))$ where $\mathcal{T}$ is a set of downstream tasks. In this paper, the proposed booster method also involves one-step gradient information into the optimization objective, aiming to attenuate the impact of harmful perturbation before fine-tuning occurs.
@@
To the best of our knowledge, we are the first to identify harmful perturbation as the cause of alignment broken, and this is our first solution that utilizes this concept to design a defense to strengthen the model’s robustness. The defense can be combined with other existing defense solutions, e.g., Vaccine (Huang et al., 2024e) and RepNoise (Rosati et al., 2024b) to further improve defense performance. A concurrent research TAR (Tamirisa et al., 2024) utilizes a similar meta-learning technique to simulate the harmful perturbation in the alignment stage. However, the insight as well as the design of our method is different from theirs. See Appendix A for a detailed discussion.
@@ 3.1 Harmful fine-tuning
Considered Scenario. Harmful fine-tuning is a security issue faced by the LLM service provider (e.g., OpenAI). The considered scenario is that users upload a few pieces of data (a subset of them are harmful) to the service provider, and the service provider fine-tune their safety-aligned foundation model with those provided datasets. The fine-tuned models are deployed in the service provider’s server and are used to serve personalized output to the users.
Threat models . We assume that the users upload a user fine-tuning dataset with $p$ (percentage) of harmful data, and other $1-p$ (percentage) of data are benign fine-tuning data. The harmful data and the benign data are considered inseparable (Qi et al., 2023; Huang et al., 2024e; Rosati et al., 2024b).
Assumptions. We assume the service provider maintains an alignment dataset (harmful prompt-safe answer pairs) and a harmful dataset (harmful prompt-harmful answer pairs) for alignment. The availability of the two pairs of data is also made in (Rosati et al., 2024b; Tamirisa et al., 2024) and both the two pairs of data are available in existing open datasets (e.g., BeaverTail (Ji et al., 2023)).
@@ 3.2 Harmful Perturbation
Next, we try to explore the concept of harmful perturbation to show the reason for alignment failure and motivate the design of our proposed defense. All the experiments are conducted by fine-tuning an aligned Llama2-7B on a mixture of SST2 and harmful data (from BeaverTails dataset). The evaluation metrics we use (e.g., harmful score) are postponed to be formally defined in Section 5.1.
@@
Definition of harmful perturbation. We refer to harmful perturbation as taking one step towards the gradient direction over the harmful data. In the fine-tuning stage, the model update rule will be $\bm{w}_{t+1}=\bm{w}_{t}-\eta g(\bm{w}_{t})$ where $g(\bm{w}_{t})$ is a stochastic gradient over a random sample of user data. If $g(\bm{w}_{t})$ is a stochastic gradient over a piece of harmful data, and we take this stochastic gradient over $\bm{w}_{t}$, this results in the so-called harmful perturbation, which should be the root reason leading to a successful harmful fine-tuning attack. We next utilize experimental results to demonstrate the impact of harmful perturbation, and how it leads to a successful attack.
@@
Impact of harmful perturbation. The left of Figure 2 shows that the model’s harmful score will substantially increase along with optimization steps invested in fine-tuning on a pure harmful dataset. On the contrary, the harmful score will not be affected much via fine-tuning on a pure SST2 dataset. This indicates that taking a step with the gradient of the harmful data (i.e., harmful perturbation) is indeed the reason for the alignment broken. We next show in the middle of Figure 2 how taking harmful perturbation and SST2 perturbation in each step will affect the harmful training loss. As shown, harmful perturbation significantly reduces the harmful training loss, which means the model starts to fit the harmful data. On the contrary, training on SST2 would only slightly increase the harmful training loss. The right of Figure 2 shows a similar trend, indicating that the fitting to the training harmful data can generalize to other unseen harmful data. In summary, our finding concludes that harmful perturbation (taking a step over the gradient of the harmful data) contributes to reduction of harmful training/testing loss, eventually triggering the model to respond in a harmful way.
@@
Figure 2: Model statistics (Left: harmful score, Middle: harmful training loss, Right: harmful testing loss) after fine-tuning on pure SST2/harmful data for different steps. Specially, harmful score measures how harmful the model is (the smaller the better), harmful training loss refers to the loss over the harmful data used in fine-tuning, while harmful testing loss refers to that over the testing harmful data that the model never sees in fine-tuning stage.
Derived Insight. As shown in our result, benign fine-tuning on SST2 will not decrease the harmful training loss, while fine-tuning on harmful data significantly decreases the harmful training loss. As a result, this paper wants to propose a method that can achieve a smaller harmful loss reduction rate when fine-tuning on the harmful data. A plausible way to achieve this goal is to modify the training objective in alignment stage such that the aligned model produced in this stage have a smaller harmful loss reduction rate when it is fine-tuned on harmful data in the later stage. From the derived insight, we further include a discussion on alternative directions in Appendix G.1.
@@ 4 Methodology
To mitigate the harmful perturbation issue, we propose an alignment stage solution to attenuate the impact of the harmful perturbation that will potentially be incurred in the fine-tuning stage. Explicitly, we want to solve the following optimization problem in the alignment stage:
@@
where $f(\bm{w})$ is the empirical loss  [^2] The empirical loss here is the cross-entropy loss if we use SFT for alignment/fine-tuning. The loss can be more complicated if RL techniques (e.g., GRPO) are adopted. over the alignment dataset and $h(\bm{w})$ is the empirical loss over the harmful dataset, $\lambda$ is the regularizer’s intensity, and $\alpha$ is the step size. Our contribution lies in the second term, which measures the gap between the original harmful loss and the harmful loss after taking a normalized step with the harmful gradient. The idea is to minimize the impact of potential harmful perturbation towards the alignment model while simultaneously minimizing its alignment loss. Specifically, $h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})$ simulates the loss after one normalized step of fine-tuning on harmful samples, and $h(\bm{w})-h(\bm{w}-\alpha\frac{\nabla h(\bm{w})}{\|\nabla h(\bm{w})\|})$ simulates the decrease of harmful loss after one step of fine-tuning on harmful samples. By minimizing this gap, the reduction rate of harmful loss after taking optimization on the real harmful samples in the fine-tuning stage will be minimized (i.e., the impact of harmful perturbation will be attenuated).
Of note, a few existing publications adopt a similar regularizer with that in Eq. (1), although they all aim to solve different problems. For example, Min et al. (2024) utilize a similar normalized gradient update regularizer (see their Eq. (3)) to make the model resistant along the backdoor-connected path. Qin et al. (2022) utilize a similar term to improve the transferability of adversarial attacks, and Zhao et al. (2022) optimize a gradient norm regularizer (see their Eq. (1)) to improve the model generalization, which can be reduced to a form similar to our regularizer in Eq. (1).
@@
where $\eta$ is the learning rate. Note that the term $\nabla(\bm{w}_{t}-\alpha\frac{\nabla h(\bm{w}_{t})}{\|\nabla h(\bm{w}_{t})\|})$ contains second-order information (i.e., Hessian Matrix), which is very computation expensive to obtain. Inspired by (Finn et al., 2017; Rajeswaran et al., 2019), we rewrite the update rule by approximating this second-order gradient term to be constant, as follows:
@@
0:  Regularizer intensity, $\lambda$; Step size, $\alpha$; Learning rate, $\eta$; SGD steps, $T$;
@@
We present the algorithm in Algorithm 1. The procedure requires three forward/backward passes for each optimization step. The first pass is to evaluate the gradient over a batch of alignment data $(\bm{x}_{t},\bm{y}_{t})$ and obtain ~$\tilde{\nabla}f(\bm{w}_{t})$. The second pass is to evaluate the harmful gradient over a batch of harmful data and obtain ~$\tilde{\nabla}h(\bm{w}_{t})$, and the third pass is to evaluate the harmful gradient again after taking a normalized harmful gradient step i.e., ~$\tilde{\nabla}h(\bm{w}_{t}-\alpha\frac{\tilde{\nabla}h(\bm{w}_{t})}{\|\tilde{\nabla}h(\bm{w}_{t})\|})$. After collecting all three gradient components, we take the final gradient step by Eq. (3).
The proposed algorithm is named Booster, named after another alignment stage solution "Vaccine" (Huang et al., 2024e). Of note, the design idea of Booster is completely different from Vaccine. In a high level, Booster is concerned with how to utilize harmful dataset to simulate the harmful perturbation, while Vaccine is concerned with how to mitigate the hidden embedding drift over the alignment dataset. Factually, the two techniques can be combined and we discuss in Appendix F.